OpenStudy (anonymous):

Find a tangent vector of unit length at the point with the given value of the parameter t. r(t) = 2 sin(t)i + 9 cos(t)j t = π/6

OpenStudy (anonymous):

find r'(pi/6), then divide the tangent vector by its length

OpenStudy (anonymous):

r'(pi/6)= 2cos(pi/6)i -9sin(pi/6) j =2(sqrt3/2) - 9(1/2) =sqrt3i - 9/2 j ?? and I am lost after this,

OpenStudy (anonymous):

sin(pi/6) = 1/2 cos(pi/6) = sqrt(3)/2 <(2)(1/2) , 9 sqrt(3)/2>= <1, (9/2)sqrt(3)> L^2 = 1^2 + [(9/2)sqrt(3)]^2 L = sqrt(61.75) (1/sqrt(61.75)) <1, (9/2)sqrt(3)>